You are analyzing the following two mutually exclusive projects and have developed the following information. What is the crossover rate? Year Project A Cash Flow Project B Cash Flow $75,000 $26,300 $29,500 $45,300 $75,000 $24,000 $26,900 $51,300 14.32 percent 13.33 percent O 15.20 percent O 14.60 percent 13.17 percent
Answer #1

Crossover rate is the rate at which the Net Present Value of both the projects is equal.

We know that NPV is equal to CashO ut flow+ Cash In flowYear1 CashI nflowYear2CashIn flowY earn (1r) 12

where n is the life of the project

r is the discounting rate

We have to determine r at which NPV of both projects is equal. So,

26,300 29,500 45 300 75,000= 24,000 26,900 51,300 75,000 +

solving both sides, we get

2, 300 2,600 -6, 000 = 0

In order to calculate r, we have to apply trial and error method.

That, is calculate NPV at two rates and then use a formula to calculate r.

So, assuming r = 15%, we get

Cash Flow Present Value Factor Present Value
1 2300 0.8696 2000
2 2600 0.7561 1965.97
3 -6000 0.6575 -3945.10
NPV 20.88

assuming r = 14%, we get

Cash Flow Present Value Factor Present Value
1 2300 0.8772 2017.54
2 2600 0.7695 2000.62
3 -6000 0.6750 -4049.83
NPV -31.67

Formula to calculate r = NPV at Lower Rate Lowerrate+ N PVat Lower Rate - NPV VatHigher Rate*(Higher Rate - Lower Rate)

ー31.67 31.67-20.88 14+ (15 14)

ー31.67 14 + 52.55

r = 14.60%

Note: Present Value Factors are calculated as

1.151 =0.8696

1.152= 0.7561

.153=0.6575

rac{1}{1.14^1}=0.8772

1.142=0.7695

rac{1}{1.14^3}= 0.6750